# Physical Intelligence

### v2

@@ Abstract
Robot learning holds tremendous promise to unlock the full potential of flexible, general, and dexterous robot systems, as well as to address some of the deepest questions in artificial intelligence. However, bringing robot learning to the level of generality required for effective real-world systems faces major obstacles in terms of data, generalization, and robustness. In this paper, we discuss how generalist robot policies (i.e., robot foundation models) can address these challenges, and how we can design effective generalist robot policies for complex and highly dexterous tasks. We propose a novel flow matching architecture built on top of a pre-trained vision-language model (VLM) to inherit Internet-scale semantic knowledge. We then discuss how this model can be trained on a large and diverse dataset from multiple dexterous robot platforms, including single-arm robots, dual-arm robots, and mobile manipulators. We evaluate our model in terms of its ability to perform tasks in zero shot after pre-training, follow language instructions from people and from a high-level VLM policy, and its ability to acquire new skills via fine-tuning. Our results cover a wide variety of tasks, such as laundry folding, table cleaning, and assembling boxes.
@@ I Introduction
Figure 2: $\bm{\pi_{0}}$ controls a mobile manipulator to fold laundry. Our model is pre-trained on diverse data from 7 distinct robot configurations and 68 tasks, and can then be used in zero-shot or fine-tuned to complex downstream tasks, as in the case of this laundry folding policy, which fetches laundry from a dryer, packs it into a hamper, brings the hamper to a folding table, and then folds each article of clothing.
@@
The contributions of our work consist of a novel generalist robot policy architecture based on VLM pre-training and flow matching, and an empirical investigation of pre-training/post-training recipes for such robot foundation models. We evaluate our model for zero-shot control with language commands, with fine-tuning to downstream tasks, and in combination with a high-level semantic policy that outputs intermediate language commands to perform complex and temporally extended tasks. While our model and system make use of a variety of ideas presented in recent work, the combination of ingredients is novel, and the empirical evaluation demonstrates a level of dexterity and generality that goes significantly beyond previously demonstrated robot foundation models. We evaluate our approach by pre-training on over 10,000 hours of robot data, and fine-tuning to a variety of dexterous tasks, including laundry folding (see Figure 2), clearing a table, putting dishes in a microwave, stacking eggs into a carton, assembling a box, and bagging groceries.
@@ II Related Work
The complexity of the tasks we illustrate goes significantly beyond prior work. While recent work has illustrated a number of more complex and dexterous behaviors, such as tying shoelaces [58] or cooking shrimp [17], we show that our framework can train very long tasks, sometimes tens of minutes in length, for behaviors that combine both physical dexterity and combinatorial complexity. For example, our laundry folding task requires the robot to manipulate a variety of clothing items that can start in any configuration, and fold multiple items in sequence. Our table bussing task requires discerning the class of novel objects (trash or dishes). We show that a single cross-embodiment model can be used as the base model for these tasks. To our knowledge, our work demonstrates the longest dexterous tasks in the end-to-end robot learning literature.
@@ IV The $\pi_{0}$ Model
where subscripts denote robot timesteps and superscripts denote flow matching timesteps, with $\tau\in[0,1]$. Recent work in high-resolution image [14] and video [38] synthesis has shown that flow matching can achieve strong empirical performance when combined with a simple linear-Gaussian (or optimal transport) probability path [28], given by $q(\mathbf{A}_{t}^{\tau}|\mathbf{A}_{t})=\mathcal{N}(\tau\mathbf{A}_{t},(1-\tau)\mathbf{I})$. In practice, the network is trained by sampling random noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, computing the “noisy actions” $\mathbf{A}_{t}^{\tau}=\tau\mathbf{A}_{t}+(1-\tau)\epsilon$, and then training the network outputs $\mathbf{v}_{\theta}(\mathbf{A}_{t}^{\tau},\mathbf{o}_{t})$ to match the denoising vector field $\mathbf{u}(\mathbf{A}_{t}^{\tau}|\mathbf{A}_{t})=\epsilon-\mathbf{A}_{t}$. The action expert uses a full bidirectional attention mask, so that all action tokens attend to each other. During training, we sample the flow matching timestep $\tau$ from a beta distribution that emphasizes lower (noisier) timesteps. See Appendix A-B for more details.
@@ VI Experimental Evaluation
Our experimental evaluation consists of zero-shot evaluation experiments that compare our base (pre-trained) model to alternative model designs, as well as detailed fine-tuning experiments that evaluate our model on challenging downstream tasks, comparing it to other methods that have been proposed for dexterous manipulation. We study the following research questions:
@@ VI-A Evaluating the base model
Figure 6: Zero-shot evaluation tasks: To evaluate our base model, we run it after pre-training on five tasks: shirt folding, bussing easy, bussing hard, grocery bagging, and toast out of toaster. The tasks require a combination of dexterous manipulation, multi-stage behaviors, and semantic recognition.
Figure 7: Zero-shot evaluation results: We evaluate $\pi_{0}$ trained for the full 700k steps, a version trained for 160k steps that matches the number of updates for baseline models, $\pi_{0}$-small, and three baselines: OpenVLA and Octo trained on all of our data, and OpenVLA trained only on the UR5e tasks (which we found to work better on UR5e tasks). Across all tasks and all comparisons, even the “parity” version of our model outperforms all baselines, and the full version of our model achieves the best results by a large margin.
@@
The evaluation metric uses a normalized score averaged over 10 episodes per task and method, where an episode receives a score of 1.0 for a full success, and a fractional score for partial success. For example, the score for bussing is the fraction of objects that are correctly placed in the proper receptacle. We describe the scoring rubrics in Appendix A-E. The results, shown in Figure 7, show that $\pi_{0}$ attains by far the best results across the board on all the zero-shot tasks, with near perfect success rates on shirt folding and the easier bussing tasks, and large improvements over all baselines. The “parity” version of $\pi_{0}$, which is trained for only 160k steps, still outperforms all the baselines, and even $\pi_{0}$-small outperforms OpenVLA and Octo. OpenVLA struggles on these tasks because its autoregressive discretization architecture does not support action chunks. The UR5e-only OpenVLA model performs better, but is still far below the performance of $\pi_{0}$. Octo does support action chunks, but has a comparatively limited representational capacity. This comparison illustrates the importance of combining large, expressive architectures with the ability to model complex distributions via flow matching or diffusion. Additionally, the comparison to $\pi_{0}$-small illustrates the importance of incorporating VLM pre-training. Unfortunately, it is hard to make this last comparison fair: $\pi_{0}$-small uses fewer parameters, but larger models are difficult to use without pre-training. Overall, these experiments show that $\pi_{0}$ provides a powerful pre-trained model with the ability to effectively perform a variety of tasks with a variety of robots, with much better performance than prior models.
@@ VI-D Mastering complex multi-stage tasks
Table bussing: This task requires bussing a table with a diverse array of novel objects in a clutter scene, presenting a much greater challenge than the benchmark in our zero-shot evaluation: the policy must generalize to unseen objects of varying shapes and sizes, and perform complex dexterous motions, such as twisting the gripper to pick up large plates and carefully grasping thin, delicate items such as glasses. The robot must handle dense clutter and intelligently sequence various behaviors — for example, to clean off a plate with trash, it must first pick up the plate, then shake its co ntents into the garbage, and then place the plate in the bin. This task is not present in pre-training.
@@
Box building: The robot has to assemble a cardboard box that starts in a flattened state. This task presents a number of major challenges: the box needs to bent in the right way, and the robot needs to hold down parts of the box while folding others, utilizing both arms and even the surface of the table to brace during folding motions. The robot might need to retry some folds, requiring a reactive and intelligent strategy. This task is not present in the pre-training data. This task is not present in pre-training.
@@
To-go box: This task requires moving several food items from a plate into a to-go box, requiring packing the items into the box so that they do not stick out, and then closing the box with both arms. This task is not present in the pre-training data. This task is not present in pre-training.
@@
Packing eggs: The robot needs to take six eggs out of a bowl and pack them into an egg carton, and then close the carton. The eggs need to be grasped in a manner appropriate to their pose inside the bowl, and then placed into open slots in the carton. This presents challenges due to the egg shape, slipperiness, and the need for careful placement. Closing the box requires the use of both arms. This task is not present in the pre-training data. This task is not present in pre-training.
@@
The results, showing average scores per task over 10 trials, are presented in Figure 13. The scoring rubrics are in Appendix A-E. A score of 1.0 represents a perfect execution, while partial scores correspond to partially completed tasks (e.g., 0.5 indicates that half the objects were bussed correctly). These tasks are very difficult, and we were not able to solve them with other methods. We therefore use these tasks to compare to ablations of our approach, evaluating $\pi_{0}$ after pre-training and fine-tuning, zero-shot evaluation after pre-training only (“zero-shot”), and training on the fine-tuning data without any pre-training (“scratch”). The results show that $\pi_{0}$ can solve many of these tasks, with our full pre-training and fine-tuning recipe performing best across the board. Note that many of these more difficult tasks show a very large improvement from using the pre-trained model, indicating that pre-training is especially useful with harder tasks. The absolute performance of $\pi_{0}$ varies across the tasks, likely due to differences in task difficulty and the degree to which the tasks are represented in pre-training. We recommend that readers watch the task videos on the accompanying website for a more complete impression of these tasks and their complexity. We believe that this level of autonomous performance on such challenging tasks represents a new state of the art in dexterous robot manipulation with learned policies.
@@ VII Discussion, Limitations, and Future Work
We presented a framework for training a robot foundation model, which we refer to as $\pi_{0}$, that consists of pre-training on highly diverse data, followed by either zero-shot evaluation or fine-tuning to complex downstream tasks. Our empirical evaluation studies tasks that combine dexterity, generalization, and temporally extended multi-stage behaviors. Our model incorporates Internet-scale vision-language model (VLM) pre-training with flow matching for representing complex high-frequency action chunks. Our pre-training mixture consists of 10,000 hours of dexterous manipulation data from 7 different robot configurations and 68 tasks, in addition to large amounts of previously collected robot manipulation data from OXE [10], DROID [23], and Bridge [52]. To our knowledge, this represents the largest pre-training mixture ever used for a robot manipulation model. Our fine-tuning experiments include over 20 tasks, where we show that our model outperforms a variety of baselines, including prior VLA models [24] and models designed specifically for dexterous manipulation [57, 9]. We also examine how our post-training recipe can enable highly complex tasks, such as folding multiple articles of clothing from arbitrary initial configurations or assembling boxes.
@@ A-B Model Architecture Details
Additional inputs and outputs. The standard PaliGemma architecture takes in a sequence of images $[\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t}]$ followed by a language prompt $\ell_{t}$. We add an input $\mathbf{q}_{t}$ for the robot’s proprioceptive state, which is mapped to the transformer embedding dimension using a linear projection. The final set of input tokens correspond to the noisy action chunk $\mathbf{A}^{\tau}_{t}=[\mathbf{a}^{\tau}_{t},...,\mathbf{a}^{\tau}_{t+H-1}]$, with the number of tokens equal to the action horizon ($H=50$ for our tasks). We only use the transformer outputs corresponding to the $H$ noisy actions, which are decoded into $\mathbf{v}_{\theta}(\mathbf{A}_{t}^{\tau},\mathbf{o}_{t})$ using a linear projection.
@@
Attention mask. $\pi_{0}$ uses a blockwise causal attention mask with 3 blocks: $[\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t},\ell_{t}]$, $[\mathbf{q}_{t}]$, and $[\mathbf{a}^{\tau}_{t},...,\mathbf{a}^{\tau}_{t+H-1}]$. Within each block, there is full bidirectional attention, whereas the tokens in each block cannot attend to the tokens in future blocks. The first block includes the input modalities from PaliGemma’s VLM pre-training, which are prevented from attending to future blocks (which include new inputs) to minimize distribution shift from said pre-training. The robot state $\mathbf{q}_{t}$ is its own block because it does not change with each flow matching integration step; preventing it from attending to the final block allows its corresponding keys and values to be cached during sampling. The final block corresponds to the noisy actions $\mathbf{A}^{\tau}_{t}$, which can attend to the full input sequence.
[…]
@@
Sampling the flow matching timestep. The original flow matching papers [28, 32] sample the flow matching timestep from a uniform distribution: $\tau\sim\mathcal{U}(0,1)$. Esser et al. [14] instead propose sampling from a logit-normal distribution that emphasizes the middle timesteps; the authors posit that at high timesteps (low noise levels), the model needs only to learn the identity function, and at low timesteps (high noise levels), the model needs only to learn the mean of the data distribution. However, we hypothesize that the task of action prediction is subtly different from high-resolution image synthesis — while it may be relatively easy to predict the mean image conditioned on a text label, predicting the mean action conditioned on a robot observation (i.e., learning $\mathbb{E}[\mathbf{A}_{t}|\mathbf{o}_{t}]$) is a much harder problem; this is because the observation $\mathbf{o}_{t}$ is very informative in that it should constrain the distribution of possible actions much more than a text label constrains the distribution of possible images. As a result, we designed a timestep sampling distribution that emphasizes low timesteps (high noise levels); additionally, timesteps above a given threshold $s$ are not sampled at all, since they are not needed so long as the integration step $\delta$ is greater than $1-s$. The distribution is given by $p(\tau)=\text{Beta}(\frac{s-\tau}{s};1.5,1)$ and is visualized in Figure 14. We use $s=0.999$ in our experiments, which allows for $\delta>\frac{1}{1000}$, or up to 1,000 integration steps.

### v4

@@ Abstract
Robot learning holds tremendous promise to unlock the full potential of flexible, general, and dexterous robot systems, as well as to address some of the deepest questions in artificial intelligence. However, bringing robot learning to the level of generality required for effective real-world systems faces major obstacles in terms of data, generalization, and robustness. In this paper, we discuss how generalist robot policies (i.e., robot foundation models) can address these challenges, and how we can design effective generalist robot policies for complex and highly dexterous tasks. We propose a novel flow matching architecture built on top of a pre-trained vision-language model (VLM) to inherit Internet-scale semantic knowledge. We then discuss how this model can be trained on a large and diverse dataset from multiple dexterous robot platforms, including single-arm robots, dual-arm robots, and mobile manipulators. We evaluate our model in terms of its ability to perform tasks via direct prompting, follow language instructions from people and from a high-level VLM policy, and its ability to acquire new skills via fine-tuning. Our results cover a wide variety of tasks, such as laundry folding, table cleaning, and assembling boxes.
@@ I Introduction
Figure 2: $\bm{\pi_{0}}$ controls a mobile manipulator to fold laundry. Our model is pre-trained on diverse data from 7 distinct robot configurations and 68 tasks, and can then either be prompted directly or fine-tuned to complex downstream tasks, as in the case of this laundry folding policy, which fetches laundry from a dryer, packs it into a hamper, brings the hamper to a folding table, and then folds each article of clothing.
@@
The contributions of our work consist of a novel generalist robot policy architecture based on VLM pre-training and flow matching, and an empirical investigation of pre-training/post-training recipes for such robot foundation models. We evaluate our model out of the box with language commands, with fine-tuning to downstream tasks, and in combination with a high-level semantic policy that outputs intermediate language commands to perform complex and temporally extended tasks. While our model and system make use of a variety of ideas presented in recent work, the combination of ingredients is novel, and the empirical evaluation demonstrates a level of dexterity and generality that goes significantly beyond previously demonstrated robot foundation models. We evaluate our approach by pre-training on over 10,000 hours of robot data, and fine-tuning to a variety of dexterous tasks, including laundry folding (see Figure 2), clearing a table, putting dishes in a microwave, stacking eggs into a carton, assembling a box, and bagging groceries.
@@ II Related Work
The complexity of the tasks we illustrate goes significantly beyond prior work. While recent work has illustrated a number of more complex and dexterous behaviors, such as tying shoelaces [58] or cooking shrimp [17], we show that our framework can learn very long tasks, sometimes tens of minutes in length, for behaviors that combine both physical dexterity and combinatorial complexity. For example, our laundry folding task requires the robot to manipulate a variety of clothing items that can start in any configuration, and fold multiple items in sequence. Our table bussing task requires discerning the class of novel objects (trash or dishes). We show that a single cross-embodiment model can be used as the base model for these tasks. To our knowledge, our work demonstrates the longest dexterous tasks in the end-to-end robot learning literature.
@@ IV The $\pi_{0}$ Model
where subscripts denote robot timesteps and superscripts denote flow matching timesteps, with $\tau\in[0,1]$. Recent work in high-resolution image [14] and video [38] synthesis has shown that flow matching can achieve strong empirical performance when combined with a simple linear-Gaussian (or optimal transport) probability path [28], given by $q(\mathbf{A}_{t}^{\tau}|\mathbf{A}_{t})=\mathcal{N}(\tau\mathbf{A}_{t},(1-\tau)\mathbf{I})$. In practice, the network is trained by sampling random noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, computing the “noisy actions” $\mathbf{A}_{t}^{\tau}=\tau\mathbf{A}_{t}+(1-\tau)\epsilon$, and then training the network outputs $\mathbf{v}_{\theta}(\mathbf{A}_{t}^{\tau},\mathbf{o}_{t})$ to match the denoising vector field $\mathbf{u}(\mathbf{A}_{t}^{\tau}|\mathbf{A}_{t})=\mathbf{A}_{t}-\epsilon$. The action expert uses a full bidirectional attention mask, so that all action tokens attend to each other. During training, we sample the flow matching timestep $\tau$ from a beta distribution that emphasizes lower (noisier) timesteps. See Appendix A-B for more details.
@@ VI Experimental Evaluation
Our experimental evaluation consists of out-of-box evaluation experiments that compare our base (pre-trained) model to alternative model designs with direct prompting, as well as detailed fine-tuning experiments that evaluate our model on challenging downstream tasks, comparing it to other methods that have been proposed for dexterous manipulation. We study the following research questions:
@@ VI-A Evaluating the base model
Figure 6: Out-of-box evaluation tasks: To evaluate our base model, we run it after pre-training on five tasks: shirt folding, bussing easy, bussing hard, grocery bagging, and toast out of toaster. The tasks require a combination of dexterous manipulation, multi-stage behaviors, and semantic recognition.
@@
Figure 7: Out-of-box evaluation results: We evaluate $\pi_{0}$ trained for the full 700k steps, a version trained for 160k steps that matches the number of updates for baseline models, $\pi_{0}$-small, and three baselines: OpenVLA and Octo trained on all of our data, and OpenVLA trained only on the UR5e tasks (which we found to work better on UR5e tasks). Across all tasks and all comparisons, even the “parity” version of our model outperforms all baselines, and the full version of our model achieves the best results by a large margin.
The evaluation metric uses a normalized score averaged over 10 episodes per task and method, where an episode receives a score of 1.0 for a full success, and a fractional score for partial success. For example, the score for bussing is the fraction of objects that are correctly placed in the proper receptacle. We describe the scoring rubrics in Appendix A-E. The results, shown in Figure 7, show that $\pi_{0}$ attains by far the best results across the board on all the out-of-box tasks, with near perfect success rates on shirt folding and the easier bussing tasks, and large improvements over all baselines. The “parity” version of $\pi_{0}$, which is trained for only 160k steps, still outperforms all the baselines, and even $\pi_{0}$-small outperforms OpenVLA and Octo. OpenVLA struggles on these tasks because its autoregressive discretization architecture does not support action chunks. The UR5e-only OpenVLA model performs better, but is still far below the performance of $\pi_{0}$. Octo does support action chunks, but has a comparatively limited representational capacity. This comparison illustrates the importance of combining large, expressive architectures with the ability to model complex distributions via flow matching or diffusion. Additionally, the comparison to $\pi_{0}$-small illustrates the importance of incorporating VLM pre-training. Unfortunately, it is hard to make this last comparison fair: $\pi_{0}$-small uses fewer parameters, but larger models are difficult to use without pre-training. Overall, these experiments show that $\pi_{0}$ provides a powerful pre-trained model with the ability to effectively perform a variety of tasks with a variety of robots, with much better performance than prior models.
@@ VI-D Mastering complex multi-stage tasks
Table bussing: This task requires bussing a table with a diverse array of novel objects in a clutter scene, presenting a much greater challenge than the benchmark in our out-of-box evaluation: the policy must generalize to unseen objects of varying shapes and sizes, and perform complex dexterous motions, such as twisting the gripper to pick up large plates and carefully grasping thin, delicate items such as glasses. The robot must handle dense clutter and intelligently sequence various behaviors — for example, to clean off a plate with trash, it must first pick up the plate, then shake its contents into the garbage, and then place the plate in the bin. This task is not present in pre-training.
@@
Box building: The robot has to assemble a cardboard box that starts in a flattened state. This task presents a number of major challenges: the box needs to bent in the right way, and the robot needs to hold down parts of the box while folding others, utilizing both arms and even the surface of the table to brace during folding motions. The robot might need to retry some folds, requiring a reactive and intelligent strategy. This task is not present in pre-training.
@@
To-go box: This task requires moving several food items from a plate into a to-go box, requiring packing the items into the box so that they do not stick out, and then closing the box with both arms. This task is not present in pre-training.
@@
Packing eggs: The robot needs to take six eggs out of a bowl and pack them into an egg carton, and then close the carton. The eggs need to be grasped in a manner appropriate to their pose inside the bowl, and then placed into open slots in the carton. This presents challenges due to the egg shape, slipperiness, and the need for careful placement. Closing the box requires the use of both arms. This task is not present in pre-training.
@@
The results, showing average scores per task over 10 trials, are presented in Figure 13. The scoring rubrics are in Appendix A-E. A score of 1.0 represents a perfect execution, while partial scores correspond to partially completed tasks (e.g., 0.5 indicates that half the objects were bussed correctly). These tasks are very difficult, and we were not able to solve them with other methods. We therefore use these tasks to compare to ablations of our approach, evaluating $\pi_{0}$ after pre-training and fine-tuning, out of the box after pre-training only (“out-of-box”), and training on the fine-tuning data without any pre-training (“scratch”). The results show that $\pi_{0}$ can solve many of these tasks, with our full pre-training and fine-tuning recipe performing best across the board. Note that many of these more difficult tasks show a very large improvement from using the pre-trained model, indicating that pre-training is especially useful with harder tasks. The absolute performance of $\pi_{0}$ varies across the tasks, likely due to differences in task difficulty and the degree to which the tasks are represented in pre-training. We recommend that readers watch the task videos on the accompanying website for a more complete impression of these tasks and their complexity. We believe that this level of autonomous performance on such challenging tasks represents a new state of the art in dexterous robot manipulation with learned policies.
@@ VII Discussion, Limitations, and Future Work
We presented a framework for training a robot foundation model, which we refer to as $\pi_{0}$, that consists of pre-training on highly diverse data, followed by either out-of-box evaluation or fine-tuning to complex downstream tasks. Our empirical evaluation studies tasks that combine dexterity, generalization, and temporally extended multi-stage behaviors. Our model incorporates Internet-scale vision-language model (VLM) pre-training with flow matching for representing complex high-frequency action chunks. Our pre-training mixture consists of 10,000 hours of dexterous manipulation data from 7 different robot configurations and 68 tasks, in addition to large amounts of previously collected robot manipulation data from OXE [10], DROID [23], and Bridge [52]. To our knowledge, this represents the largest pre-training mixture ever used for a robot manipulation model. Our fine-tuning experiments include over 20 tasks, where we show that our model outperforms a variety of baselines, including prior VLA models [24] and models designed specifically for dexterous manipulation [57, 9]. We also examine how our post-training recipe can enable highly complex tasks, such as folding multiple articles of clothing from arbitrary initial configurations or assembling boxes.
@@ A-B Model Architecture Details
Additional inputs and outputs. The standard PaliGemma architecture takes in a sequence of images $\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t}]$ followed by a language prompt $\ell_{t}$. We add an input $\mathbf{q}_{t}$ for the robot’s proprioceptive state, which is mapped to the transformer embedding dimension using a linear projection. The final set of input tokens correspond to the noisy action chunk $\mathbf{A}^{\tau}_{t}=[\mathbf{a}^{\tau}_{t},...,\mathbf{a}^{\tau}_{t+H-1}]$, with the number of tokens equal to the action horizon ($H=50$ for our tasks). We only use the transformer outputs corresponding to the $H$ noisy actions, which are decoded into $\mathbf{v}_{\theta}(\mathbf{A}_{t}^{\tau},\mathbf{o}_{t})$ using a linear projection.
@@
Attention mask. $\pi_{0}$ uses a blockwise causal attention mask with 3 blocks: $\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t},\ell_{t}]$, $\mathbf{q}_{t}]$, and $\mathbf{a}^{\tau}_{t},...,\mathbf{a}^{\tau}_{t+H-1}]$. Within each block, there is full bidirectional attention, whereas the tokens in each block cannot attend to the tokens in future blocks. The first block includes the input modalities from PaliGemma’s VLM pre-training, which are prevented from attending to future blocks (which include new inputs) to minimize distribution shift from said pre-training. The robot state $\mathbf{q}_{t}$ is its own block because it does not change with each flow matching integration step; preventing it from attending to the final block allows its corresponding keys and values to be cached during sampling. The final block corresponds to the noisy actions $\mathbf{A}^{\tau}_{t}$, which can attend to the full input sequence.
Action expert. $\pi_{0}$ is implemented as a single transformer with two sets of weights (also known as experts [45]), where each token is routed to one of the experts; the weights interact only through the transformer’s self-attention layers. The images and language prompt, $\mathbf{I}^{1}_{t},...,\mathbf{I}^{n}_{t},\ell_{t}]$, are routed to the larger VLM backbone, which we initialize from PaliGemma. The inputs not seen during VLM pre-training, $\mathbf{q}_{t},\mathbf{A}^{\tau}_{t}]$, are routed to the action expert. PaliGemma is based on the Gemma 2B [49] language model, which uses multi-query attention [44] and a configuration of {width=2048, depth=18, mlp_dim=16,384, num_heads=18, num_kv_heads=1, head_dim=256}. Since the experts interact only in the self-attention layers, width and mlp_dim do not necessarily need to match between experts. To speed up inference (which requires multiple forward passes of the action expert), we downsize the action expert to {width=1024, mlp_dim=4096}, resulting in a parameter count of $\sim$300M.
@@
Sampling the flow matching timestep. The original flow matching papers [28, 32] sample the flow matching timestep from a uniform distribution: $\tau\sim\mathcal{U}(0,1)$. Esser et al. [14] instead propose sampling from a logit-normal distribution that emphasizes the middle timesteps; the authors posit that at high timesteps (low noise levels), the model needs only to learn the identity function, and at low timesteps (high noise levels), the model needs only to learn the mean of the data distribution. However, we hypothesize that the task of action prediction is subtly different from high-resolution image synthesis — while it may be relatively easy to predict the mean image conditioned on a text label, predicting the mean action conditioned on a robot observation (i.e., learning $\mathbb{E}[\mathbf{A}_{t}|\mathbf{o}_{t}]$) is a much harder problem; this is because the observation $\mathbf{o}_{t}$ is very informative in that it should constrain the distribution of possible actions much more than a text label constrains the distribution of possible images. As a result, we designed a timestep sampling distribution that emphasizes low timesteps (high noise levels); additionally, timesteps above a given threshold $s$ are not sampled at all, since they are not needed so long as the integration step $\delta$ is greater than 1s1-s. The distribution is given by $p(\tau)=\text{Beta}(\frac{s-\tau}{s};1.5,1)$ and is visualized in Figure 14. We use $s=0.999$ in our experiments, which allows for $\delta>\frac{1}{1000}$, or up to 1,000 integration steps.